OpenStudy (anonymous):

Given the function f(x) = 5x - 9, find f(-3).

OpenStudy (anonymous):

plug in -3 into the function

jimthompson5910 (jim_thompson5910):

replace each copy of x with -3, then evaluate f(x) = 5x - 9 f(-3) = 5(-3) - 9 f(-3) = ???

OpenStudy (anonymous):

It gives you -3 for x, so plug in -3 to the equation. 5 (-3) - 9 -15 - 9 = 24